# Single-Shot Readout and Weak Measurement of a Tin-Vacancy Qubit in Diamond

### v1

@@ IV.1 Controlled dephasing
Aside from the practical use of characterizing efficiency and optimizing readout, Fig. 4 illustrates that the power dependence of readout offers a broad pathway for exploring quantum measurement. Measurement is fundamental to quantum theory, and is taken as one of the axioms of quantum mechanics [47]. Under the action of measurement, a quantum state undergoes probabilistic collapse into one of its eigenstates. This collapse destroys quantum superposition, i.e. causes decoherence.
After a projective measurement, for example as demonstrated in Fig. 3, a measured quantum state is entirely dephased and is projected into a classical, probabilistic mixture of eigenstates. In contrast, for the case of a weak measurement [51], a quantum state has a finite chance of collapse and therefore, on average, retains some coherence. This equivalently limits the classical information a detector may acquire about its state. In this section, we combine the techniques we have thus-far demonstrated: high-fidelity spin control (Fig. 2), single-shot projective measurement (Fig. 3) and weak measurement (Fig. 4) in order to study measurement induced dephasing of the ${}^{-}$’s electronic spin, Fig. 5. This study is an interesting parallel to recent work in the superconducting qubit community, e.g. [51, 52, 53, 54, 55], and contrasts from previous color-center based work focused on nuclear spin control [56, 57, 58, 59].
To study measurement induced dephasing we do the experiment described in Fig. 5: the qubit is prepared in a superposition state with initial coherence $|\hat{\rho}_{\downarrow\uparrow}|$, subjected to a weak measurement, and finally, projectively measured to determine its remaining post-measurement coherence, $|\hat{\rho}_{\downarrow\uparrow}^{\prime}|$. The weak measurement pulse consists of a resonant laser drive of power $p$ and duration $\tau$, and is placed at the center of a dynamical decoupling sequence (see Fig. 11d for details). As $\tau$ and $p$ increase we observe a decrease in readout contrast, which is proportional to the qubit’s remaining coherence after undergoing weak measurement. We fit this to the exponential model $|\hat{\rho}_{\downarrow\uparrow}^{\prime}|=|\hat{\rho}_{\downarrow\uparrow}|e^{-\Gamma_{\phi}\tau}$, where $\Gamma_{\phi}$ is the measurement induced dephasing rate.
@@
where $R$ is given by Eq. 2. Notice there is a simple relationship $\Gamma_{\phi}\approx\chi\kappa/2$ between Eq. 1 and Eq. 7. Polarization, dephasing, and emission are all closely related with power dependence governed by the relative ratio of pump power to its saturation value, $p/(p+p_{\mathrm{sat}})$, as follows from the optical Bloch equations, and with the overall timescale of dynamics set by the optical decay rate $\gamma$.
This experiment offers a way to precisely study how strongly the laser drive interacts with a qubit. In the low power limit, dephasing rises linearly with $p$. Above saturation, dephasing must plateau to the limit controlled by the optical decay rate, $\gamma$. A fit of Eq. 7 to the data in Fig. 5d determines a saturation power of $p_{\mathrm{sat}}=1436\pm 21\,\mathrm{nW}$, specified at the input to the cryostat. We attribute the difference between $p_{\mathrm{sat}}$ here compared to the measurement in Fig. 4 to drift in the optical setup between datasets, and/or, that the data in Fig. 4 is post-selected on CRCs whereas data in Fig. 5 is not.
@@ IV.2 A general method to characterize efficiency
Apart from the fundamental scientific curiosity surrounding study of quantum measurement, examining the low power limits of measurements depicted in both Fig. 3c and Fig. 5d offers an alternative means to evaluate measurement efficiency. Far below saturation, both the mean photon count difference $\bar{n}_{b}-\bar{n}_{d}$ and the measurement induced dephasing rate $\Gamma_{\phi}$ scale linearly with power. Taylor expanding Eq. 6 in the $p/p_{\mathrm{sat}}\ll 1$ limit and for $\delta=0$ gives: $\bar{n}_{b}-\bar{n}_{d}=Ap+\mathcal{O}(p^{2})$ with proportionality constant $A=\gamma\eta\tau/(2p_{\mathrm{sat}})=1.744\times 10^{-2}\pm 8.305\times 10^{-5}\,\mathrm{nW^{-1}}$. At low power, $\Gamma_{\phi}=Bp+\mathcal{O}(p^{2})$ is also linear with proportionality constant $B=\gamma/(4p_{\mathrm{sat}})=41.5\pm 2.4\,\mathrm{kHz/nW}$. Therefore,
@@
where $A$ is a fit of Eq. 6 to the weak measurement data in Fig. 4b, and $B$ is a fit of Eq. 7 to the measurement induced dephasing data in Fig. 5e, and $\tau=50\,\mathrm{\mu s}$ is the readout integration window. See Ref. [52] for a similar approach for characterizing the efficiency of superconducting qubit measurement.
@@
Despite this difference, we include Eq. 8 because it provides a general framework that may be useful in future experiments requiring precise characterization of solid-state atomic systems. This measurement allows for determination of efficiency within the low power limit, thus avoiding the effect of power-dependent qubit instability. It may be particularly applicable to other color centers which are difficult to saturate or have a competing ionization process.
@@ V Conclusion
In conclusion, we demonstrate single-shot readout of a single electronic spin of an ${}^{-}$ qubit in diamond, along with coherent control of this spin by a microwave drive. We report a readout fidelity of $F_{r}=87.4\%$ using a $50\,\mathrm{\mu s}$ readout integration window. These results are achieved by aligning the magnetic field near the spin dipole direction (misalignment of $\varphi\approx 10\degree$) to operate at cyclicity $\chi\approx 2200$. To our knowledge this is highest fidelity single-shot readout of an ${}^{-}$ spin that has been published to-date: the readout in Ref. [39] reports a lower fidelity of $74\%$, and uses an un-strained emitter for which microwave spin control is not feasible. (Ref. [32] reports readout of an ${}^{-}$ nuclear spin, and also does not include spin control.) Next, using high-fidelity readout we repeatedly measure the qubit and characterize the extent to which subsequent measurements are correlated; doing so we characterize a conditional fidelity of $F_{c}=98.5\%$ (Eq. 4), and a QND-equivalent fidelity of $F_{q}\approx 77\%$. All of these results are obtained operating at a measurement efficiency of only $\eta\approx 0.1\%$, indicating orders-of-magnitude room for improvement in future devices that utilize nanophotonics. Finally, we use rapid microwave spin control and long spin coherence to study dephasing induced by measurement. This demonstrates the fundamental interplay between measurement and dephasing that is inherent to quantum mechanics. It is also a versatile technique to characterize how strongly a laser drive interacts with a qubit, and to characterize measurement efficiency.
@@
Our work advances the ${}^{-}$ as a platform for building quantum networks. Taken together with other recent ${}^{-}$ work including nanophotonic integration [29, 30, 34, 33], microwave spin control [26, 27], spin-photon entanglement [31], and nuclear spin states [36, 32], the ${}^{-}$ has all the features necessary for building a quantum repeater node of similar scale and complexity to ${}^{-}$ based devices [15, 16, 17], but which favorably are robust to heating effects and elevated temperatures. More broadly, the understanding and methods we develop here – in particular the use of weak quantum measurement and measurement induced dephasing – serve as powerful metrological tools and are applicable to the study of a wide variety of solid-state atom-like systems.
@@
Acknowledgments.— This work has been supported by the Department of Energy under the Q-NEXT program, and Grant No. DE-SC0020115 and by the National Science Foundation Award ECCS 2150633. E.I.R. and C.P.A. acknowledge support by an appointment to the Intelligence Community Postdoctoral Research Fellowship Program at Stanford University administered by Oak Ridge Institute for Science and Education (ORISE) through an interagency agreement between the U.S. Department of Energy and the Office of the Director of National Intelligence (ODNI). J.C. acknowledges support from the Terman Faculty Fellowship at Stanford. H.C.K. acknowledges support by the Burt and Deedee McMurtry Stanford Graduate Fellowship (SGF). J.G. acknowledges support from the Hertz Fellowship. G.S. and S.A. acknowledge support from the Stanford Bloch Postdoctoral Fellowship. D.R. acknowledges support from the Swiss National Science Foundation (Project No. P400P2_194424). D.R. and S.A. contributed to this work prior to joining AWS. This work was performed, in part, at the Center for Integrated Nanotechnologies, an Office of Science User Facility operated for the U.S. Department of Energy (DOE) Office of Science. Sandia National Laboratories is a multimission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International, Inc., for the U.S. DOE’s National Nuclear Security Administration under contract DE-NA-0003525. The views expressed in the article do not necessarily represent the views of the U.S. DOE or the United States Government. We thank Jacob Feder, Jonathan Marcks, Mia Froehling Gallier and Cyrus Zeledon for help with instrument control code, based on the “nspyre” framework [60]. We thank Haiyu Lu, Shuo Li, Patrick McQuade, Zhi-Xun Shen and Nicholas Melosh for assistance with diamond sample preparation. We thank Nazar Delgan and F. Joseph Heremans for collaboration on the preparation of related devices. Part of this work was performed at the Stanford Nanofabrication Facility (SNF) and the Stanford Nano Shared Facilities (SNSF), supported by the NSF under Grant No. ECCS-2026822.
@@ A.1.1 Three-level model
In this section we model the dynamics of an ${}^{-}$ center under a resonant drive. This will be used to model spin polarization, photon emission during readout, and measurement induced dephasing. Consider a three-level system (a “$\Lambda$-system”) consisting of two ground states $\ket{\downarrow}$ and $\ket{\uparrow}$, with an excited state |$\ket{A}$, Fig. 6. This is a simplified version of the energy diagram in Fig. 2a. Relaxation from |$\ket{A}$ occurs with probability $P_{\mathrm{A1}}$ or $P_{\mathrm{A2}}$ into the $\ket{\downarrow}$ or $\ket{\uparrow}$ states, respectively, and cyclicity $\chi=P_{\mathrm{A1}}/P_{\mathrm{A2}}$. Since $P_{\mathrm{A1}}+P_{\mathrm{A2}}=1$ we have $P_{\mathrm{A2}}=1/(1+\chi)$ and $P_{\mathrm{A1}}=\chi/(1+\chi)$.
@@
Under a Rabi drive the Hamiltonian in a rotating frame is:
@@
where the first term contains the optical Rabi rate $\Omega_{\mathrm{A1}}$, and the second term describes its detuning $\delta$ from the A1 transition. We also define stochastic quantum jump operators:
@@
with $\gamma_{\mathrm{A1}}=\gamma\chi/(1+\chi)$ and $\gamma_{\mathrm{A2}}=\gamma/(1+\chi)$, and where $\gamma=\gamma_{\mathrm{A1}}+\gamma_{\mathrm{A2}}$ is the optical decay rate given by the lifetime of the |$\ket{A}$ state.
@@ A.1.2 Two-level model
only. Here, rates $R_{1}=RP_{\mathrm{A1}}=R\chi/(1+\chi)$ and $R_{2}=RP_{\mathrm{A2}}=R/(1+\chi)$ are:
@@
$\ \rho_{\downarrow\downarrow}[t]=|\alpha|^{2}e^{-\kappa t},$ (17)
@@
$\ \rho_{\uparrow\uparrow}[t]=1+(|\beta|^{2}-1)e^{-\kappa t},$ (20)
where $\Gamma_{\phi}=R/2$ and $\kappa=R/(1+\chi)$. The qubit therefore undergoes drive-induced dephasing at rate $\Gamma_{\phi}$, and polarization at the slower rate $\kappa$.

### v2

@@ IV.1 Controlled dephasing
Aside from the practical use of characterizing efficiency and optimizing readout, Fig. 4 illustrates that the power dependence of readout offers a broad pathway for exploring quantum measurement. Measurement is fundamental to quantum theory and is taken as one of the axioms of quantum mechanics [52]. Under the action of measurement, a quantum state undergoes probabilistic collapse into one of its eigenstates. This collapse destroys quantum superposition, i.e., causes decoherence.
After a projective measurement, for example, as demonstrated in Fig. 3, a measured quantum state is entirely dephased and is projected into a classical, probabilistic mixture of eigenstates. In contrast, for the case of a weak measurement [58], a quantum state has a finite chance of collapse and therefore, on average, retains some coherence. This equivalently limits the classical information a detector may acquire about its state. In this section, we combine the techniques we have thus far demonstrated: high-fidelity spin control (Fig. 2), single-shot projective measurement (Fig. 3), and weak measurement (Fig. 4) in order to study measurement-induced dephasing of the SnV-’s electronic spin (Fig. 5). This study is an interesting parallel to recent work in the superconducting qubit community, e.g. [58, 59, 60, 61, 62], and contrasts from previous color center work focused on nuclear spin control [63, 64, 65, 66]. This study also contrasts from cavity QED-based projective readout of color center electronic spins [19, 20, 21] but, importantly, has a direct analogy to the measurement-induced decoherence of a nuclear spin state in a cavity-coupled SiV-, studied in Ref. [22].
To study measurement-induced dephasing we do the experiment described in Fig. 5: The qubit is prepared in a superposition state with initial coherence $|\hat{\rho}_{\downarrow\uparrow}|$, subjected to a weak measurement, and finally, projectively measured to determine its remaining postmeasurement coherence, $|\hat{\rho}_{\downarrow\uparrow}^{\prime}|$. The weak measurement pulse consists of a resonant laser drive of power $p$ and duration $\tau$ and is placed at the center of a dynamical decoupling sequence (see Fig. 12d for details). As $\tau$ and $p$ increase we observe a decrease in readout contrast, which is proportional to the qubit’s remaining coherence after undergoing weak measurement. We fit this to the exponential model $|\hat{\rho}_{\downarrow\uparrow}^{\prime}|=|\hat{\rho}_{\downarrow\uparrow}|e^{-\Gamma_{\phi}\tau}$, where $\Gamma_{\phi}$ is the measurement-induced dephasing rate.
@@
where $R$ is given by Eq. 2. Notice there is a simple relationship $\Gamma_{\phi}\approx\Lambda\Gamma_{p}/2$ between Eqs. 1 and 7. Polarization, dephasing, and emission are all closely related with power dependence governed by the relative ratio of pump power to its saturation value, $p/(p+p_{\mathrm{sat}})$, as follows from the optical Bloch equations, and with the overall timescale of dynamics set by the optical decay rate $\gamma$.
This experiment offers a way to precisely study how strongly the laser drive interacts with a qubit. In the low-power limit, dephasing rises linearly with $p$. Above saturation, dephasing must plateau to the limit controlled by the optical decay rate $\gamma$. A fit of Eq. 7 to the data in Fig. 5d determines a saturation power of $p_{\mathrm{sat}}=1436\pm 21\,\mathrm{nW}$, specified at the input to the cryostat. We attribute the difference between $p_{\mathrm{sat}}$ here compared to the measurement in Fig. 4 to drift in the optical setup between datasets and/or that the data in Fig. 4 are post-selected on CRCs whereas data in Fig. 5 are not.
We note that the CRC pass fraction does not change with power until $p\gg p_{\mathrm{sat}}$ (see Fig. 15b), and we furthermore note that an off-resonant control pulse does not cause any measurement-induced dephasing (Fig. 18). These experiments indicate that the observed rise in dephasing rate with drive power is not due to an alternate effect such as drive-induced blinking or spectral diffusion, a conclusion which is consistent with prior demonstrations of coherent optical Raman spin control [29, 30, 31].
@@ IV.2 A general method to characterize efficiency
IV.2.1 Photoluminescent readout
Apart from the fundamental scientific curiosity surrounding study of quantum measurement, examining the low-power limits of measurements depicted in both Fig. 3c and 5e offers an alternative means to evaluate measurement efficiency. Far below saturation, both the mean photon count difference $\bar{n}_{b}-\bar{n}_{d}$ and the measurement-induced dephasing rate $\Gamma_{\phi}$ scale linearly with power. Taylor expanding Eq. 6 in the $p/p_{\mathrm{sat}}\ll 1$ limit and for $\delta=0$ gives $\bar{n}_{b}-\bar{n}_{d}=Ap+\mathcal{O}(p^{2})$ with proportionality constant $A=\gamma\eta\tau/(2p_{\mathrm{sat}})=1.744\times 10^{-2}\pm 8.305\times 10^{-5}\,\mathrm{nW^{-1}}$. At low power, $\Gamma_{\phi}=Bp+\mathcal{O}(p^{2})$ is also linear with proportionality constant $B=\gamma/(4p_{\mathrm{sat}})=41.5\pm 2.4\,\mathrm{kHz/nW}$. Therefore,
@@ IV.2.1 Photoluminescent readout
where $A$ is a fit of Eq. 6 to the weak measurement data in Fig. 4b, $B$ is a fit of Eq. 7 to the measurement-induced dephasing data in Fig. 5e, and $\tau=50\,\mathrm{\mu s}$ is the readout integration window. See Ref. [59] for a similar approach for characterizing the efficiency of superconducting qubit measurement.
@@
Despite this difference, we include Eq. 8 because it provides a general framework that may be useful in future experiments requiring precise characterization of solid-state atomic systems. This measurement allows for determination of efficiency within the low-power limit, thus avoiding the effect of power-dependent qubit instability. It may be particularly applicable to other color centers which are difficult to saturate or have a competing ionization process.
IV.2.2 Dispersive readout
The fundamental physics of weak quantum measurement is agnostic to technique, platform, or type of detector. The introduced framework for accurately characterizing the measurement efficiency also applies to other types of readout, including the case of dispersive measurement of cavity QED systems. Dispersive readout is an important ingredient to establishing quantum network nodes based on coherent atom-light interactions [20, 19, 21, 22, 23].
In dispersive readout, discrimination between the $\ket{\uparrow}$ and $\ket{\downarrow}$ states is governed by the coupling between the spin state and cavity photons, causing spin-dependent scattering of a probe tone. Weak and strong measurement regimes can also be understood systematically based on the same characterization of measurement-induced dephasing used in Fig. 5, such that a quantitative comparison between the dephasing rate $\Gamma_{\phi}$ and the collected information about the qubit state (e.g., signal $\bar{n}_{b}-\bar{n}_{d}$) may be used to determine measurement efficiency $\eta$. See Appendix A.4 for a detailed mathematical discussion.
As with the photoluminescence-based readout discussed elsewhere in this work, efficiency is important for dispersive readout, because greater efficiency allows for high-fidelity measurement using lower probe powers and shorter readout durations, thus mitigating drive-induced sources of error. We also expect that precise characterization of efficiency will be an important tool for the development of quantum repeater nodes, in parallel to its general application to benchmarking sophisticated circuit QED systems [59, 62]. After all, careful measurement of loss is a first step toward its mitigation, and reducing loss will improve the entanglement generation rate within a quantum network.
@@ V Conclusion
In conclusion, we demonstrate single-shot readout of a single electronic spin of an SnV- qubit in diamond, along with coherent control of this spin by a microwave drive. We report a readout fidelity of $F_{r}=87.4\%$ using a $50\,\mathrm{\mu s}$ readout integration window. These results are achieved by aligning the magnetic field near the spin dipole direction (misalignment of $\varphi\approx 10\degree$) to operate at cyclicity $\Lambda\approx 2200$. To our knowledge, this is highest fidelity single-shot readout of an SnV- spin that has been published to-date: The readout in Ref. [43] reports a lower fidelity of $74\%$, and uses an unstrained emitter for which microwave spin control is not feasible. (Reference [36] reports readout of an SnV- nuclear spin and also does not include spin control.) Next, using high-fidelity readout, we repeatedly measure the qubit and characterize the extent to which subsequent measurements are correlated; doing so, we characterize a conditional fidelity of $F_{c}=98.5\%$ (Eq. 4), and a QND-equivalent fidelity of $F_{q}\approx 77\%$. All of these results are obtained operating at a measurement efficiency of only $\eta\approx 0.1\%$, indicating orders-of-magnitude room for improvement in future devices that utilize nanophotonics for higher collection efficiency.
Finally, we use rapid microwave spin control and long spin coherence to study dephasing induced by measurement. This demonstrates the fundamental interplay between measurement and dephasing that is inherent to quantum mechanics. It is also a versatile technique to characterize how strongly a laser drive interacts with a qubit and to characterize measurement efficiency. As an immediate next step, it will be of interest to apply the readout techniques developed in this manuscript to dispersive readout of cavity QED systems. In direct analogy with its use in circuit QED [59, 60, 61, 62, 67], studying the spin-dependent interaction of a probe tone with a cavity will be a useful way to benchmark measurement efficiency and understand sources of loss.
@@
Overall, our work advances the SnV- as a platform for building quantum networks. Taken together with other recent SnV- work including nanophotonic integration [33, 34, 38, 37], microwave spin control [30, 31], spin-photon entanglement [35], and nuclear spin states [40, 36], the SnV- has all the features necessary for building a quantum repeater node of similar scale and complexity to SiV--based devices [19, 20, 21, 22] but which favorably are robust to heating effects and elevated temperatures. More broadly, the understanding and methods we develop here— in particular, the use of weak quantum measurement and measurement-induced dephasing—serve as powerful metrological tools and are applicable to the study of a wide variety of solid-state atomlike systems.
While in press, the authors became aware of relevant work in Ref. [68].
@@
Acknowledgments.— This work has been supported by the Department of Energy under the Q-NEXT program and Grant No. DE-SC0020115 and by the National Science Foundation Grant No. ECCS 2150633. E.I.R. and C.P.A. acknowledge support by an appointment to the Intelligence Community Postdoctoral Research Fellowship Program at Stanford University administered by Oak Ridge Institute for Science and Education (ORISE) through an interagency agreement between the U.S. Department of Energy and the Office of the Director of National Intelligence (ODNI). J.C. acknowledges support from the Terman Faculty Fellowship at Stanford. H.C.K. acknowledges support by the Burt and Deedee McMurtry Stanford Graduate Fellowship (SGF). J.G. acknowledges support from the Hertz Fellowship. G.S. and S.A. acknowledge support from the Stanford Bloch Postdoctoral Fellowship. D.R. acknowledges support from the Swiss National Science Foundation (Project No. P400P2_194424). D.R. and S.A. contributed to this work prior to joining AWS. This work was performed, in part, at the Center for Integrated Nanotechnologies, an Office of Science User Facility operated for the U.S. Department of Energy (DOE) Office of Science. Sandia National Laboratories is a multimission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International, Inc., for the U.S. DOE’s National Nuclear Security Administration under Contract No. DE-NA-0003525.
The views expressed in the article do not necessarily represent the views of the U.S. DOE or the United States Government. We thank Jacob Feder, Jonathan Marcks, Mia Froehling Gallier, and Cyrus Zeledon for help with instrument control code, based on the “nspyre” framework [69]. We thank Haiyu Lu, Shuo Li, Patrick McQuade, Zhi-Xun Shen and Nicholas Melosh for assistance with diamond sample preparation. We thank Nazar Delgan and F. Joseph Heremans for collaboration on the preparation of related devices. We thank Noah Lordi for helpful discussions. Part of this work was performed at the Stanford Nanofabrication Facility (SNF) and the Stanford Nano Shared Facilities (SNSF), supported by the NSF under Grant No. ECCS-2026822.
@@ A.1.1 Three-level model
In this section, we model the dynamics of an SnV- center under a resonant drive. This is used to model spin polarization, photon emission during readout, and measurement-induced dephasing. Consider a three-level system (a “lambda system”) consisting of two ground states $\ket{\downarrow}$ and $\ket{\uparrow}$, with an excited state |$\ket{A}$, Fig. 6. This is a simplified version of the energy diagram in Fig. 2a. Relaxation from |$\ket{A}$ occurs with probability $P_{\mathrm{A1}}$ or $P_{\mathrm{A2}}$ into the $\ket{\downarrow}$ or $\ket{\uparrow}$ states, respectively, and cyclicity $\Lambda=P_{\mathrm{A1}}/P_{\mathrm{A2}}$. Since $P_{\mathrm{A1}}+P_{\mathrm{A2}}=1$ we have $P_{\mathrm{A2}}=1/(1+\Lambda)$ and $P_{\mathrm{A1}}=\Lambda/(1+\Lambda)$.
@@
Under a Rabi drive the Hamiltonian in a rotating frame is
@@
where the first term contains the optical Rabi frequency $\Omega_{\mathrm{A1}}$, and the second term describes its detuning $\delta$ from the A1 transition. We also define stochastic quantum jump operators:
@@
with $\gamma_{\mathrm{A1}}=\gamma\Lambda/(1+\Lambda)$ and $\gamma_{\mathrm{A2}}=\gamma/(1+\Lambda)$, and where $\gamma=\gamma_{\mathrm{A1}}+\gamma_{\mathrm{A2}}$ is the optical decay rate given by the lifetime of the |$\ket{A}$ state.
@@ A.1.2 Two-level model
only. Here, rates $R_{1}=RP_{\mathrm{A1}}=R\Lambda/(1+\Lambda)$ and $R_{2}=RP_{\mathrm{A2}}=R/(1+\Lambda)$ are
@@
$\ \rho_{\downarrow\downarrow}[t]=|\alpha|^{2}e^{-\Gamma_{p}t},$ (17)
@@
$\ \rho_{\uparrow\uparrow}[t]=1+(|\beta|^{2}-1)e^{-\Gamma_{p}t},$ (20)
where $\Gamma_{\phi}=R/2$ and $\Gamma_{p}=R/(1+\Lambda)$. The qubit, therefore, undergoes drive-induced dephasing at rate $\Gamma_{\phi}$, and polarization at the slower rate $\Gamma_{p}$.